Fermi National Laboratory


Accelerator Update for January 23, 2006 - January 25, 2006

Monday January 23

The day shift began with Tevatron (Tev) A4 cooldown in progress, the Recycler Ring (RR) stash at 129.3 E10, Antiproton Source (Pbar) stack at 16.5 E10, and with MiniBooNE, NuMI and MCenter (MIPP) taking beam.

Pbar studies continue

At 10:17 AM, MIPP was disabled for shots from the RR to the Main Injector (MI).

At 1:21 PM, MIPP is off because of a power supply problem with S:HP3US.

S:HP3US is repaired and beam returns to MIPP at 3:26 PM.

At 5:08 PM, operations began a search and secure of the Tevatron.

Tev secure was completed by 9:08 PM.

Operations began D0 search and secure at 10:07 PM.

Tuesday January 24

The midnight shift began with the Tevatron cold and turning on, the Recycler Ring (RR) stash at 16.2 E10, RR stash is at 0 due to studies, and with MiniBooNE, NuMI and MIPP taking beam.

Tevatron is ramping at 2:20 AM.

Pbar was turned off at 4:00 AM to prepare for 10 hour access period on Tuesday.

Booster and MI were turned off at 5:00 AM to prepare for access period.

PBar, Booster and MI accesses began at 6:00 AM.

D0 and CDF went into access at 8:25 AM.

Transfer Hall and NuMI also went into access at 8:57 AM.

MiniBoone begins access at 9:27 AM.

CDF is out of access and secure by 12:15 PM.

Linac was turned on at 1:47 PM

At 6:19 PM a vacuum leak found in the AP2 line will delay stacking in Pbar.

Beam in Booster at 6:32 PM.

At 8:07 PM, the Tevatron was ramping.

Beam to the MI at 8:48 PM.

Beam to MiniBooNe at 9:01 PM.

Beam to NuMI at 9:27 PM.

D0 access was completed at 9:46 PM.

Beam to MIPP at 10:16 PM.

Some beam to the Tev at 10:51 PM.

At 11:02 PM, problems with power supply D:V730 further delay Pbar stacking.

Wednesday January 25

The midnight shift began with Tevatron studies, the Recycler Ring (RR) stash at 12.3 E10, Pbar down with vacuum and power supply problems, and with MiniBooNE, NuMI and MIPP taking beam.

At 3:02 AM, the Tev began having problems with the TEL. Experts were called in to investigate.

Pbar power supply D:V730 was repaired by 6:45 AM.

At 7:07 AM, experts are accessing the transfer hall to replace a bad modulator tube for the TEL.

At 7:09 AM, Pbar is still unable to turn on because power is out at F23, Duty Electrician is investigating.

The Plan for Wednesday

Try to resume stacking in Pbar, then transfer stack to the RR, continue with Tev startup, run beam to NuMI, MiniBooNE and Switchyard.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 01/25/2006    email Fermilab